OpenStudy (anonymous):

10x^3-640x=0 solve by factoring

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=10x%5E3-640x%3D0 x=-8 x=0 x=8

OpenStudy (anonymous):

did this help

OpenStudy (aum):

10x^3 - 640x = 0 10x(x^2 - 64) = 0 10x(x+8)(x-8) = 0 x = 0 or x = -8 or x = 8.

OpenStudy (anonymous):

good job aum